You have a wire of length L = 1.6 m for making a square coil of a dc motor. The current in the coil is I = 1.1 A, and the magnetic field of the motor has a magnitude of B = 0.14 T. Find the maximum torque exerted on the coil when the wire is used to make (a) a single-turn square coil and (b) a two-turn square coil.
Answer #1

Solution)

Given, length l= 1.6 m

Current, I= 1.1 A

Magnetic Field, B= 0.14 T

We know, maximum torque, T= NIAB

Where, no of turns, N=1

Part a) For single coil

We have,

Torque= NI*(L/4)^2*B = 1*(1.1)*(1.6/4)^2*0.14 =0.02464 Nm

=====

Part b)

For two turn, we have N=2

So, torque= NI*(L/8)^2*B

Torque= 2* 1.1 * (1.6/8)^2*0.14 = 0.01232 Nm (Ans)

======